The product of two consecutive integers is 29 more than the sum. What are the integers?
Answer by mananth(16949)

let the integers be x,x+1
x(x+1)= x+x+1+29
x^2+x=2x+1+29
x^2-2x+x-30=0
x^2-x-30=0
(x-6)(x+5)=0
x= 6 or -5
6,7
-5,-4
...
m.ananth@hotmail.ca